A methods and measurements analyst needs to develop a time standard for a certain task. In a preliminary study, he observed one of his workers perform this task five times, with the following results:

Observation 1 2 3 4 5
Time (secs) 81 79 75 83 84

What is the standard time for this task if the employee worked at a 29 percent faster pace than average and an allowance of 16 percent of the workday is used?

(Keep one decimal place in your answer)
Answer #1

Average of given time,

Observed time = 81+79+75+83+84 / 5 = 80.4

Normal time = Observed time * Performance rating = 80.4 * 1.29 = 103.716

Allowance factor = 1 / 1 - 0.16 = 1.19

Standard time = 103.716 * 1.19 = 123.4 seconds.

**** PLEASE LEAVE A LIKE. THANK YOU.****